• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name cusanno

    Log entry time 16:03:08 on March 1,2009

    Entry number 264086

    keyword=Day Shift Summary

    8:05 failure in Left HRS Q1, cooling level < 15%, After few minutes the cooling is back, but Q1 is too "hot" and reset fails due to overtemperature. We wait for a while to reset it again

    8:20 beam line trip, the same happened during the owl shift, MCC says

    8:25 beam back

    8:27 Left Q1 still failing, Gary paged. He says that if remote reboot fails we need an access to reboot manually

    8:37 Asked for controlled access

    8:45 Brad called saying about the hardware reset on the middle room, it works, we cancel the controlled access

    9:00 Q1 recovered, we go back into production

    9:24 beam trips again, still problem in vacuum line

    9:29 beam back

    9:32 BB channel 14.6 tripped

    9:45 beam trips again

    10:35 Beam is coming back

    12:55 NMR measurement

    13:35 Beam off for a while

    14:05 Beam is back

    15:00 We changed kinematics into G, reference cell in

    15:15 calibration run with N2 reference cell

    15:55 He-3 target moved in

    16:00 start production at kinematics G